1) 13+(-3)^2+4(-3)+1-[-10-(-6)]
---------------------------------------
{[4+5]/[4^2-3^2(4-3)-8]}+12

Answer by CubeyThePenguin(3113)

numerator: 13 + (-3)^2 + 4(-3) + 1 - (-10 - (-6))
= 13 + 9 - 12 + 1 + 4
= 15

denominator: [(4 + 5)/(4^2 - 3^2(4 - 3) - 8)] + 12
= (9/(16 - 9 - 8)) + 12
= 9/(-1) + 12
= 3

fraction: 15/3 = 5